### v1

@@ Abstract
The remainder of this paper is organised as follows. Section 2 presents an overview of other publicly available datasets for benthic mapping. Section 3 and Section 4 describe the details of the surveys conducted for acoustic and optical data acquisition, including the system setup and sensor characteristics, and outline how the collected data was processed. Section 5, on the other hand, presents details on the additional processing of the collected data to compile datasets for training ML models. Section 6 explains how the datasets are structured. Section LABEL:sec:benchmarking evaluates the performance of standard segmentation models retrained on the collected SSS dataset for seafloor classification. Finally, Section 7 presents conclusions and directions for future work.
@@ 5.2 Multi-modal Dataset
The geographical coordinates of the acoustic and optical data were recorded using different reference frames. The former, collected on the tow-fish, utilizes East-North-Up (ENU) coordinates, while the latter, acquired onboard the Girona 1000 AUV, employs North-East-Down (NED) coordinates. Due to this difference, when computing the intersection between the two sets of footprints, the optical data was first converted to ENU format.
@@
where and denote the easting and northing coordinates at ping , and are the rotations about the and axes at ping , is the altitude at ping , is the number of bins inside ping , and is the slant resolution. The rotation about the horizontal axis helps determine the width and the offset geographical coordinates of the SSS ping, since it will cause the SSS pulses to travel further away.
@@ 6 Repository Structure
The dataset is provided in raw and processed formats so as not to constrain users to a specific application. All data have been synchronised and timestamped to make it easily accessible. The data is hosted and available at https://doi.org/10.7910/DVN/EXFFWP. Exemplar code and detailed instructions on how to process the data are available at https://github.com/CIRS-Girona/opti-acoustic-preprocessing-tools.
The dataset is organised into two main directories: raw and datasets. The raw directory contains all the unprocessed files, including ROS Bag recordings (together with the associated calibration metadata) from the optical imaging surveys, XTF files from the SSS data, and .xyz files from the MBES data. The SSS and MBES folders are further subdivided into sector-specific subdirectories, each containing the corresponding raw data files for that sector. Additionally, the GIS folder contains the georeferenced mosaics and ArcGIS interpretations. Figure 9 presents an overview of the repository structure.
/ raw camera survey-0.bag survey-1.bag survey-2.bag … calibration.json SSS N01 transect-0.xtf transect-1.xtf … N02 N03 … MBES N01 transect-0.xyz transect-1.xyz … N02 N03 … GIS datasets SSS pre-training fine-tuning SSS-CAM
The datasets directory contains the SSS and opti-acoustic datasets. Each of the SSS datasets is further divided into train and test subfolders. The test set consists of independent transects that do not overlap with those in the train set. Note that the pre-training dataset consists of SSS tiles per sector, whereas the fine-tuning dataset consists of annotated waterfalls rather than tiles, in order to avoid constraining users. A Python script to cut the waterfalls into overlapping tiles has been provided in the GitHub repository. The opti-acoustic dataset, on the other hand, comprises two folders: sonar and camera, along with the three metadata files: sonar.csv, camera.csv and correspondences.json. The sonar and camera folders store all the matched SSS tiles and optical images. This is as presented in Figure 10.
datasets SSS-CAM sonar tile-0.png tile-1.png … sonar.csv camera image-0.png image-1.png image-2.png … camera.csv correspondences.json

### v2

@@ Abstract
The remainder of this paper is organised as follows. Section 2 presents an overview of other publicly available datasets for benthic mapping. Section 3 and Section 4 describe the details of the surveys conducted for acoustic and optical data acquisition, including the system setup and sensor characteristics, and outline how the collected data was processed. Section 5, on the other hand, presents details on the additional processing of the collected data to compile datasets for training ML models. Section 6 explains how the datasets are structured. Finally, Section 7 presents conclusions and directions for future work.
@@ 5.2 Multi-modal Dataset
where and denote the easting and northing coordinates at ping , and are the yaw and pitch rotations at ping , is the altitude at ping , is the number of bins inside ping , and is the slant resolution. The pitch angle helps determine the width and the offset geographical coordinates of the SSS ping, since a pitching movement will cause the SSS pulses to travel further away.
@@ 6 Repository Structure
The data is provided in both raw and processed formats so as not to constrain users to a specific application. The data has also been synchronised and timestamped to make it easily accessible. The data is available at https://dataverse.harvard.edu/dataverse/benthicat. Exemplar code and detailed instructions on how to process the data are available at https://github.com/CIRS-Girona/opti-acoustic-preprocessing-tools.
The data is organised as a collection of individual datasets. The collection currently comprises four datasets:
BenthiCat - Raw, which contains all the unprocessed files, including ROS Bag recordings (together with the associated calibration metadata) from the optical imaging surveys, XTF files from the SSS data, and .xyz files from the MBES data. The SSS and MBES folders are further subdivided into sector-specific subdirectories, each containing the corresponding raw data files for that sector. Additionally, the GIS folder contains the georeferenced mosaics and ArcGIS interpretations. Figure 9 presents an overview of the repository structure.
/ camera survey-0.bag survey-1.bag survey-2.bag … calibration.json SSS N01 transect-0.xtf transect-1.xtf … N02 N03 … MBES N01 transect-0.xyz transect-1.xyz … N02 N03 … GIS
Figure 9: Structure of the raw dataset. -
BenthiCat - SSS Pre-training that contains unlabelled SSS tiles for unsupervised representation learning and self-supervised pre-training tasks.
BenthiCat - SSS Seafloor Segmentation containing the manually annotated SSS waterfalls for supervised fine-tuning. The dataset is divided into train and test subsets, with the test set consisting of independent transects that do not spatially overlap with those in the train set.
The dataset contains waterfalls instead of pre-cut tiles to avoid constraining users to a fixed tile size; however, a Python script to generate overlapping tiles from the waterfalls is available in the GitHub repository.
BenthiCat - SSS-CAM, which contains the spatially matched optical and SSS samples. It includes two folders: sonar and camera, along with the three metadata files: sonar.csv, camera.csv and correspondences.json. The sonar and camera folders store all the matched SSS tiles and optical images, respectively. An overview of the repository structure is presented in Figure 10.
/ sonar tile-0.png tile-1.png … sonar.csv camera image-0.png image-1.png image-2.png … camera.csv correspondences.json
Figure 10: Structure of the opti-acoustic dataset.